OpenStudy (niahloves):

What is the product? -5 x (-3) x (-7) x (2) x (0.5) A. –210 B. –105 C. 105 D. 210

OpenStudy (niahloves):

@ltrout

OpenStudy (ltrout):

For future reference, use * instead of x to multiply. You'll learn why when you get into algebra. But Let's look at your problem. There's an uneven amount of negative numbers, so that means the final result will be negative. Now you start multiplying. Any order is fine. -5 * -3 = 15. -7 * 2 = -14. 15 * -14 = -210 -210 * .5 (or 1/2 of -210) = -105.

OpenStudy (niahloves):

One more please

OpenStudy (niahloves):

I'm thinking its C,or D

OpenStudy (ltrout):

Easy way to tell. Find any one with uneven amounts of negative numbers. a = 4 negative. b = 2 negative. c = 3 negative*. d = 3 negative, but it's multiplying by a 0 so the end result would be 0, which is not negative. So C would be the correct one.

OpenStudy (niahloves):

I was correct i picked c :)

OpenStudy (niahloves):

I mastered my test,Thank you!!!!!